A postal truck leaves its station and heads to Chicago ,averaging 40 mph. An error in the mailing schedule is spotted and 24 minutes after the truck leaves, a car is sent to overtake the truck. If the car averages 50mph, how long will it take to catch the postal truck?

Answer by checkley77(12844)

D=RT
D=40X
D=50(X-.4)
SEEING AS THE DISTANCES ARE THE SAME SET THEM EQUAL TO EACH OTHER.
40x=50(x-.4)
40x=50x-20
40x-50x=-20
-10x-20
x=-20/-10
x=2 HOURS THEY WILL MEET.
40*2=50(2-.4)
80=50*1.6
80=80